late the pH of a mixture that is 0.095 M in HF and 0.225 M in HC6H5O.
Answer #1

mixture of 0.095M HF and O.da5M (64504 dissociate Both HF and Pa and na Consoh 6hSon are wea are weak acid ka of 4F = 7.2810| 2 HF | 0.075 H+ +F- o o н x € 0.015-x о tx x to x ka= [ht] [F] [IF ] (x) (x) (0.015-2) 7.2410-4 = x <co.0157-24104 ? 0.095 6.84x10% = x² 6.843785 X 0.0083 M = x X = 0.0083 [F] = [4+] = 0.0083 gholt (64500 Now consider dissociatin C6